# Heterogeneous Treatment Effects and

### v1

@@ F.2 Assumptions on the Latent Utility Distributions
In practice, utility is usually specified as a linear function $V^{i}_{m}=X\beta$, where $X$ is the vector of observed variables and $\beta$, are parameters. We can estimate $\beta$ from the data. Then, we could treat $V_{m}^{i}$ as the directly-affected outcome to explore the mechanism through HTE. It should be clear that Propositions 1 and 2 hold for $V_{i}^{m}$ even if the real utility is $U$.
Now, we use our motivating example to illustrate how to apply the above RUM and distribution assumptions in practice. In the main text, we provide a simple model of the effect of exogenous shocks on pro-incumbent voting. The model specifies the systematic component of utility, $V_{m}^{i}$. A RUM requires us to specify a random component of utility, $\varepsilon$ which also affects individual behavior and is additively separable from $V_{m}^{i}$ (i.e. not relevant to the theory of interests). Therefore, given the observed voting data we have, if we assume the distribution of $\varepsilon$ is type-I extreme value, then the probability individual votes for the incumbent is
@@
Appendix G Simulation
Illustration: The distinction between directly-affected and indirectly-affected outcomes is novel to this paper. To illustrate the logic and implications of learning about mechanisms from HTE in the case of an indirectly-affected outcome, we provide a short simulation that incorporates real attitudinal data. Specifically, we consider a hypothetical persuasion experiment that aims to shift support for greenhouse gas regulation among partisans in the US. Consistent with approaches used by scholars of persuasion, we will examine heterogeneity in partisan affiliation (here, simplified to Democrats and Republicans) (see Coppock, 2022: etc.). We use data on (1) partisan affiliation; (2) support for greenhouse gas regulation, coded as a binary outcome where $1$ designates support for increased regulation; and (3) demographic covariates from the 2020 American National Election Study. It is useful to note that partisans’ opinions are relatively polarized on this issue: while 82.2% (95% CI: [80.5%, 84.0%]) of Democrats favor increasing regulations, just 38.3% (95% CI: [35.7%, 40.9%]) of Republicans favor such regulations. This suggests that partisanship is strongly prognostic of support for greenhouse gas regulation.
Various theories of learning or additudinal change incorporate mechanisms that imply that partisans may react differently to information about greenhouse gas regulation. Little, Schnakenberg, and Turner (2022) classify two mechanisms for belief formation and attitude change: accuracy and directional motives. Within their model, ideology (partisanship) is posited as a moderator of directional motives but not accuracy motives, meaning that partisanship is a candidate MIV for directional motives. To this end, we simulate different processes of attitudinal change to examine when we observe HTE in partisanship. Our simulation proceeds as follows:
@@ Appendix G Simulation
We vary $\tau\in[-1.5,1.5]$, which are treatment effects on a logistic scale. [^22] The assumption of a constant $\tau$ is clearly a simplification for the purposes of illustration. It does not generally follow from the Little, Schnakenberg, and Turner (2022) model. Figure A.6 reports the results of our simulation. In the left panel, we see that for non-zero treatment effects (e.g., for any $\tau\neq 0$), we always observe HTE in partisanship, even when effects on the latent scale are homogeneous, e.g., the degree of attitudinal change is not moderated by partisanship. We observe different treatment effects for Democrats and Republicans on the binary outcome even with homogeneous treatment effects on the latent attitude because of different densities of respondents about the relevant cutpoint in the latent variable (see Figure A.7).

### v2

@@ G.2 Assumptions on the Latent Utility Distributions
In practice, utility is usually specified as a linear function $V^{i}_{m}=X\beta$, where $X$ is the vector of observed variables and $\beta$, are parameters. We can estimate $\beta$ from the data. Then, we could treat $V_{m}^{i}$ as the directly-affected outcome to explore the mechanism using HTE. It should be clear that Propositions 1 and 2 hold for $V_{i}^{m}$ even if the utility is $U$.
Now, we use our motivating example to illustrate how to apply the above RUM and distribution assumptions in practice. In the main text, we provide a simple model of the effect of exogenous shocks on pro-incumbent voting. The model specifies the systematic component of utility, $V_{m}^{i}$. A RUM requires us to specify a random component of utility, $\varepsilon$ which also affects individual behavior and is additively separable from $V_{m}^{i}$ (i.e. not relevant to the theory of interest). Therefore, given the observed voting data we have, if we assume the distribution of $\varepsilon$ is type-I extreme value, then the probability individual votes for the incumbent is
@@
Appendix H Simulation
Illustration: The distinction between directly-affected and indirectly-affected outcomes is novel to this paper. To illustrate the logic and implications of learning about mechanisms from HTEs in the case of an indirectly-affected outcome, we provide a short simulation that incorporates real attitudinal data. Specifically, we consider a hypothetical persuasion experiment that aims to shift support for greenhouse gas regulation among partisans in the US. Consistent with approaches used by scholars of persuasion, we will examine heterogeneity in partisan affiliation (here, simplified to Democrats and Republicans) (see Coppock, 2022: etc.). We use data on (1) partisan affiliation; (2) support for greenhouse gas regulation, coded as a binary outcome where $1$ designates support for increased regulation; and (3) demographic covariates from the 2020 American National Election Study. It is useful to note that partisans’ opinions are relatively polarized on this issue: while 82.2% (95% CI: [80.5%, 84.0%]) of Democrats favor increasing regulations, just 38.3% (95% CI: [35.7%, 40.9%]) of Republicans favor such regulations. This suggests that partisanship is strongly prognostic of support for greenhouse gas regulation.
Various theories of learning or additudinal change incorporate mechanisms that imply that partisans may react differently to information about greenhouse gas regulation. Little, Schnakenberg, and Turner (2022) classify two mechanisms for belief formation and attitude change: accuracy and directional motives. Within their model, ideology (partisanship) is posited as a moderator of directional motives but not accuracy motives, meaning that partisanship is a candidate MDV for directional motives. To this end, we simulate different processes of attitudinal change to examine when we observe HTEs in partisanship. Our simulation proceeds as follows:
@@ Appendix H Simulation
We vary $\tau\in[-1.5,1.5]$, which are treatment effects on a logistic scale. [^19] The assumption of a constant $\tau$ is clearly a simplification for the purposes of illustration. It does not generally follow from the Little, Schnakenberg, and Turner (2022) model. Figure A.7 reports the results of our simulation. In the left panel, we see that for non-zero treatment effects (e.g., for any $\tau\neq 0$), we always observe HTEs in partisanship, even when effects on the latent scale are homogeneous, e.g., the degree of attitudinal change is not moderated by partisanship. We observe different treatment effects for Democrats and Republicans on the binary outcome even with homogeneous treatment effects on the latent attitude because of different densities of respondents about the relevant cutpoint in the latent variable (see Figure A.8).